6. A person consumes two goods, bacon (B) and tofu (T). They get utility from consumption in the following way U(B,T)-4B2T. The person has $50 to spend on the two goods and plans to spend all $50. Bacon costs $1 and tofu costs $2. Set up your objective function of the form (YOU DO NOT HAVE TO SOLVE, JUST SET IT UP): Max U such that budget constraint is satisfied.
Answer #1

6. The utility function is given as U =4B-T . The income is $50, and price of B is $1 and T is $2. The budget constraint would be 2T + B = 50 . The maximization problem (in terms of objective function and constraint) would be as below.

Mar U = 4B-T

such that 2T + B = 50 .

______________________________________________________________________________________

Note : The solution of the problem can be done via Lagrange's multiplier method as below.

The Lagrangian function of the maximization problem is L = 4B-T + (50 - 2T - B) . The FOCs are as below.

aL or (4B-T + (50 - 27 - B)) = 0 or 8BT + \lambda (- 1) = 0 or 8BT = \lambda .

aL aT or (4 BºT + (50 – 2T - B)) = 0 or 4B + X(-2)=0 or 4B^2 = 2\lambda or P=ad .

aL or (4B2T + X(50 – 2T - B)) = 0 or 2T + B = 50 .

Solving for the first two FOCs, we have 8BT = 2B^2 or 4T = B , which is the utility maximizing combination of products.

Putting this in the budget constraint, we have 2T + (4T) = 50 or T^* = 25/3 , and as B^* = 4T^* , we have B^* = 100/3 . These are the optimal combination of goods which would maximize the utility given the budget constraint.